# A Methodological Report on Anomaly Detection on Dynamic Knowledge Graphs

### v1

@@ 1 Introduction
Dynamic Knowledge Graphs (DKGs), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time. By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as social media [23], the Internet of Things (IoT) [22, 21], and e-commerce [^2] https://eval.ai/web/challenges/challenge-page/2267/overview.
On the leading Microservices Architecture (MSA) platform, Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms. These data are typically stored in DKGs in the TTL format. Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of errors and anomalies.
@@
To tackle these challenges, this paper proposes a new approach to anomaly detection on dynamic knowledge graphs using graph structure and temporal dependencies. This approach combines neighbourhood feature extraction with support vector machines (SVMs), extreme gradient boosting (XGB), and self-attention mechanisms. Our approach significantly outperforms the baseline performance on the ’Dynamic Knowledge Graph Anomaly Detection’ challenge dataset at ISWC 2024, providing a robust solution for anomaly detection in dynamic, complex environments.
The structure of the paper is as follows. Section 2 gives the formulation of the problem and describes related work. Section 3 presents the methodology of the paper, while section 4 describes the experimental part of the work and discusses the results obtained. Section 5 concludes our contribution to the ADDKG challenge and discusses future work.
@@ 2.1 Dynamic Knowledge Graphs
Combined with the dynamic graph representations, the DKG is a heterogeneous dynamic graph represented by timestamped edge lists. Among the DKG representation formats, TTL (Turtle) is commonly used. A typical TTL allows for the serialisation of triples, which can include nodes, edges, and their attributes, along with timestamps, making TTL versatile for representing entities and relations in an MSA system. An example of a TTL-based DKG is given in Eq.1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
$\text{DKG}=\{(s,p,o,t)\mid s,o\in\text{Nodes},p\in\text{Predicate},t\in\text{Timestamp}\}$ (1)
In the problem studied in this paper, anomalous data is represented by its entity $s$, the time interval during which the anomaly occurs $(t_{start},t_{end})$, and the class of the anomaly $c$.
@@ 2.2 Anomaly Detection
Traditional anomaly detection methods typically rely on statistical techniques and distance-based measures. However, these methods often struggle with high-dimensional data and complex dependencies [27]. To overcome these problems, common machine learning approaches include support vector machines (SVM) [6], isolation forests (IF) [7], and extreme gradient boosting (XGB) [8]. SVM [6] identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers. Isolation Forest [7] is an unsupervised method and detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness. XGB [8] is a gradient boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in distinguishing normal from abnormal data.
With the advancement of software algorithms and hardware devices, deep learning methods have been increasingly applied to anomaly detection due to their ability to extract complex features [16]. Common architectures for time series include recurrent neural networks (RNN) [9], temporal convolutional networks (TCN) [12], and the self-attention mechanism (SA) [13]. RNN (Recurrent Neural Network) is a type of neural network designed to process sequential data by capturing temporal dependencies. LSTM (Long Short-Term Memory) [10] and GRU (Gated Recurrent Unit) [11] are variants of RNNs designed to solve the vanishing and exploding gradient problems in long sequence training. TCN (Temporal Convolutional Network) [12] captures long-term temporal dependencies with causal convolutions. Self-attention mechanisms capture long-range dependencies by computing the relationships between all elements in a sequence, making them particularly effective for identifying complex patterns. For graph anomaly detection, Graph Neural Networks (GNNs) propagate information through graph structures to identify anomalous patterns [19, 18]. Despite the aforementioned studies proposing numerous algorithms for anomaly detection, significant challenges remain when applied to dynamic knowledge graphs.
2.3 Challenges of Anomaly Detection on Dynamic Knowledge Graph
Due to the complex temporal and structural dependencies in dynamic knowledge graphs, traditional machine learning methods struggle to capture the underlying features. However, deep learning models typically require large amounts of labelled data for effective training, and anomalous data is inherently scarce.
Although anomaly detection tasks on dynamic graphs without attributed edges can generate anomalous edges through methods such as negative sampling [19, 18, 20], this approach becomes infeasible in the case of dynamic knowledge graphs with attributed edges. Therefore, effectively capturing the temporal and structural dependencies of features on dynamic knowledge graphs is crucial to tackling this challenge.
@@ 3 Methodology
This section first introduces three representations for dynamic knowledge graphs studied in this task: sequential data, one-hop graph structure, and two-hop graph structure, with each representation incorporating increasingly complex structural information. We then describe the machine learning/deep learning methods applied to each representation. At the end of the section, we propose ensemble schemes for combining multiple methods.
@@ 3.1 DKG as Sequential Data
The time-evolving attributes of nodes can be obtained by extracting raw TTL data. For example, in MSA, common node categories could be $\mathcal{C}$ representing "Cluster", $\mathcal{P}$ representing "Pod", $\mathcal{S}$ representing "Service", and $\mathcal{N}$ representing "Node". For each given node category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{N_{k}\times T\times d_{k}}$, where $N_{k}$ indicated the number of instances of the category $k$, $T$ indicates the number of snapshots, and $d_{k}$ indicates the number of features.
@@
Therefore, the sequential data representation of a given category in the DKG can be expressed by a matrix $\mathbf{F}\in\mathbb{R}^{T\times D_{k}}$ as shown in Eq. 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
$\mathbf{F}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
Once the matrix representation of the attributes $\mathbf{F}$ is obtained, additional information can be calculated. For example, the difference captures the trend of the time series, while the variance reflects the stability of the data fluctuations. Specific formulas for statistical properties are given in Eq. 4.
$\text{Difference}(t)$ $=$ $Z(t){-}Z(t{-}1)$ (4)
$\text{Variance}(t)$ $=$ $\frac{1}{\tau}\sum_{j=t-\tau+1}^{t}(\mathbf{Z}(j)-\bar{\mathbf{Z}})^{2}$
$\text{Rolling Mean}_{\tau}(t)$ $=$ $\frac{1}{\tau}\sum_{j=t-\tau+1}^{t}\mathbf{Z}(j)$
$\text{Rolling Sum}_{\tau}(t)$ $=$ $\sum_{j=t-\tau+1}^{t}\mathbf{Z}(j)$
@@ 3.2 One-Hop Graph Structure
To explore the structural information of the knowledge graph, we propose to use one-hop neighbour aggregation to capture the relationships between different entities. Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig.1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp based on the temporal information considered in the previous subsection.
@@
For instance, to encode and predict data for the category Service, information related to the Service can be aggregated. Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, and Connections can be aggregated to form the attributes of the Service, as shown in Eq. 5 where $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the set of Node instances related to $P_{k}$. In this manner, multiple one-hop graph structures are used for message propagation, and these structural-temporal features provide the model with a richer context.
Figure 1: The dataset consists of knowledge graphs representing an e-commerce Kubernetes cluster at specific timestamps (UTC). Each snapshot provides information about the state of the microservice platform, deployments, resources, and component relationships. The graphs are in TTL format and detail subjects, objects, and their relationships based on a defined ontology. The ontology overview illustrates the relationships between Kubernetes concepts.
@@ 3.3 Two-Hop Graph Structure
Some graph machine learning algorithms have pointed out that a larger convolutional field, or aggregation of multi-hop neighbours, can optimise the model. Inspired by this, we realised that aggregating second-hop neighbours could provide richer information. In the tree-like graph of the MSA problem, considering only the current level and its associated root level forms a bipartite graph. For nodes of category $k$, the two-hop neighbours on this bipartite graph are still nodes of the same category.
Recognising that there may be interdependencies between Service instances at each timestamp, we propose to use two-hop neighbour aggregation. Intuitively, this aggregation can be achieved by concatenating vertex features, as shown in Eq. 6 where $N_{\mathcal{S}}$ represents the number of Service instances and $S_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.
$\mathbf{X}_{\text{service}}(t)=[\mathbf{S_{1}}(t),\mathbf{S_{2}}(t),\ldots,\mathbf{S_{N_{S}}}(t)];$ (6)
Another method of capturing two-hop neighbours is using the self-attention mechanism to capture the correlations among vertices. The self-attention mechanism works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others. This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data.
@@ 3.4 Ensemble Learning
Ensemble learning is a prevalent methodology in the domain of machine learning that mitigates variance through the consolidation of multiple models, thereby enhancing model robustness. The integration of diverse and superior-performing models is a prerequisite for the efficacy of ensemble learning [14]. Ensemble learning allows multiple models to make predictions on the same problem and thus vote on the final prediction. The voting principle can be hard voting or soft voting, while the voting mechanism usually consists of unanimous voting and majority voting, which will be introduced in this subsection.
@@
As two common voting mechanisms combined with hard voting, especially in anomaly detection, unanimous voting classifies an instance as an anomaly only if all the models agree. Majority voting, on the other hand, classifies an instance as an anomaly if the majority of models agree.
@@ 4.1 Experimental Setup
All experiments were conducted utilising the ISWC 2024 ADDKG dataset. The dataset contains five days of labelled data, with three days of data used for training and the remaining two days used for testing. The data was collected at approximately 15 second intervals and presents a challenge due to its imbalance, with anomalies representing only approximately 4% of the total data. To mitigate this imbalance, class weights were applied to the loss functions during model training to ensure that the models effectively learned to detect anomalies.
As outlined in the methodology section, our approach included three levels of representation. For the first representation, sequential modelling is employed. The original training set is randomly divided into two subsets: 80% for training, and the other 20% for validation. For the other two representations, a more sophisticated approach to data splitting is under consideration. For normal class, the 80-20 split mechanism remains applicable. However, in the case of the anomaly class, it is necessary to identify complete and continuous sequences of anomalous records as "anomalous events". The objective was to maintain the anomalous events in their entirety during the process of data splitting, the splitting of a single anomalous process between the training and validation sets. This approach ensures the contextual integrity of the anomalies, allowing the models to learn more effectively from complete anomalous events. To achieve this, for each anomaly type, one complete event was randomly selected for validation, while the remaining events were used for training.

### v2

@@ 1 Introduction
Dynamic Knowledge Graphs (DKG), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time. By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as social media [23], the Internet of Things (IoT) [22, 21], and e-commerce [^2] https://eval.ai/web/challenges/challenge-page/2267/overview.
On the leading Micro-Services Architecture (MSA) platform, such as Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms. These data are typically stored in DKGs in the TTL format [^3] https://en.wikipedia.org/wiki/Turtle_(syntax). Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of anomalies.
@@
To address these challenges, this paper introduces a novel anomaly detection framework for dynamic knowledge graphs, leveraging both graph structure and temporal dependencies. Our method integrates neighbourhood feature extraction with advanced machine learning techniques, including Support Vector Machine (SVM) [6], Extreme Gradient Boosting (XGB) [8], and Self-Attention mechanism (SA) [13] to enhance the model’s ability to capture complex patterns of anomalies. Our approach achieves significant performance improvements over the baseline in the Anomaly Detection on Dynamic Knowledge Graphs challenge (ADDKG) [^4] https://eval.ai/web/challenges/challenge-page/2267/overview at ISWC 2024, providing a robust solution for anomaly detection in dynamic and complex environments.
The structure of the paper is as follows. Section 2 gives the formulation of the problem and describes related work. Section 3 presents the methodology of the paper. Section 4 describes the experimental part of the work and discusses the results obtained. Section 5 concludes our contribution to the ADDKG challenge and discusses future work.
@@ 2.1 Dynamic Knowledge Graphs
Combined with the dynamic graph representations, the Dynamic Knowledge Graph (DKG) is modelled as a heterogeneous dynamic graph, represented by timestamped edge lists. Among the various formats for representing DKGs, Turtle (TTL) is one of the most commonly used. TTL allows the serialisation of nodes, edges, their attributes and timestamps, making it a versatile format for representing entities and relationships in a multi-agent system (MSA). An example of a TTL-based DKG is given in Eq. 1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
$\text{DKG}=\{(s,p,o,t)\mid t\in[t_{1},t_{2},...,T]\}$ (1)
In the problem studied in this paper, anomalous data is represented by its entity $s$, the time interval during which the anomaly occurs $(t_{start},t_{end})$, and the class of the anomaly $c$. Table 1 shows all the anomaly types $c$ in the dataset.
Table 1: Description of system faults and their types.
Fault Type Description
Node CPU Hog Static Consumes the CPU resources of the targeted node.
Node Drain Dynamic Forces rescheduling of the resources running on
the targeted node.
Node Memory Hog Static Consumes the memory resources of the targeted
node.
Pod Autoscaler Dynamic Changes the desired number of replicas for the
targeted workload.
Pod CPU Hog Static Consumes the CPU resources of the targeted
application container.
Pod Delete Dynamic Causes forced/graceful pod failure of a random
replica of the targeted workload
Pod Network Latency Static Latency introduced in the network communication
between different pods.
@@ 2.2 Anomaly Detection
Traditional anomaly detection methods typically rely on statistical techniques and distance-based measures. However, these methods often struggle with high-dimensional data and complex dependencies [27]. To overcome these problems, common machine learning approaches include Support Vector Machine (SVM) [6], eXtreme Gradient Boosting (XGB) [8], and Isolation Forest (IF) [7]. SVM identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers. XGB is a gradient-boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in separating normal from abnormal data. Isolation Forest is an unsupervised method that detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
With the advancement of software algorithms and hardware, deep learning methods have been increasingly popular for anomaly detection due to their ability to extract complex features [16]. Common architectures for time series include Recurrent Neural Networks (RNN) [9], Temporal Convolutional Networks (TCN) [12], and the Self-Attention mechanism (SA) [13]. RNNs, including Long Short-Term Memory (LSTM) [10] and Gated Recurrent Unit (GRU) [11], capture temporal dependencies, addressing gradient issues in long sequences. TCNs handle long-term dependencies through causal convolutions, while SA captures long-range dependencies by examining relationships across sequences. For graph anomaly detection, Graph Neural Networks (GNN) propagate information through graph structures to identify anomalous patterns [19, 18].
Despite advancements in anomaly detection algorithms, significant challenges persist in dynamic knowledge graphs. Traditional machine learning methods struggle to capture the complex temporal and structural dependencies, while deep learning models require large amounts of labelled data, which is often scarce. Although techniques like negative sampling [19, 18, 20] can generate anomalous edges in dynamic graphs without attributed edges, this approach is infeasible for dynamic knowledge graphs with attributed edges. Capturing both temporal and structural dependencies remains key to tackling this challenge.
@@ 3 Methodology
This section introduces three representations of dynamic knowledge graphs: sequential data, one-hop, and two-hop graph structures, each adding increasing structural complexity. We then describe the machine learning and deep learning methods applied to these representations, followed by proposed ensemble learning schemes. Figures 1 and 2 provide an overview of the pipeline, including the data structure, the three levels of the data processing, and the final output.
Figure 1: Left: The ontology overview of the relationships between Kubernetes concepts of the ADDKG dataset. Right: The simplified hierarchical structure of the ADDKG dataset. The container information is aggregated as the attributes of Pods. The Connections serve as attributes of the edges between Pods and Services.
Figure 2: This figure presents the methodology used in this study, with three levels of processing: sequential data, one-hop, and two-hop graph structures. Each level introduces increasing structural complexity, showing how these representations are fed into anomaly detection models.
@@ 3.1 DKG as Sequential Data
The time-evolving attributes of nodes can be obtained by extracting raw TTL data, without considering the structural information of the DKG. As shown in the "DKG as Sequential Data" part of Fig. 2, we merge all the sequential data of all the vertices in a DKG by the timestamp. This includes vertices of type $\mathcal{C}$ representing "Cluster", type $\mathcal{P}$ representing "Pod", type $\mathcal{S}$ representing "Service", and type $\mathcal{N}$ representing "Node". For each given node category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{N_{k}\times T\times d_{k}}$, where $N_{k}$ indicated the number of instances of the category $k$, $T$ indicates the number of snapshots, and $d_{k}$ indicates the number of features. A sliding window then captures temporal dependencies by concatenating the historical information of the past $\tau$ time steps, as shown in Eq. 2.
@@
Therefore, the sequential data representation of a given category $k$ in the DKG can be expressed by a matrix $\mathbf{X}\in\mathbb{R}^{T\times D_{k}}$, where $D_{k}=N_{k}\times d_{k}$, as shown in Eq. 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
$\mathbf{X}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
Once the matrix representation of the attributes $\mathbf{X}$ is obtained, additional information can be computed. For example, the difference captures the trend of the time series, while the variance reflects the stability of the data fluctuations.
@@ 3.2 One-Hop Graph Structure
To explore the structural information of the knowledge graph, we propose to use one-hop neighbour aggregation to capture the relationships between different entities. Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig. 1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp, as shown the aggregate method in the "One-Hop Chart Structure" section of Fig. 2.
@@
For instance, to encode and predict data for the category Service, information related to the Service can be aggregated. Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, Containers and Connections can be aggregated to form the attributes of the Service, as shown in Eq. 4, where $n_{S}$ represents the number of Service instances, $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the elements in the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the elements in the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the elements in the set of Node instances related to $P_{k}$. In this manner, multiple one-hop graph structures are used for message propagation, and these structural-temporal features provide the model with a richer context.
@@ 3.3 Two-Hop Graph Structure
Some graph machine learning algorithms have pointed out that a larger convolutional field, or aggregation of multi-hop neighbours, can optimise the model. Inspired by the potential of aggregating second-hop neighbours to provide richer information, we observed that in the tree-like graph of the MSA problem, considering only the current level and its root forms a bipartite graph. For nodes of category $k$, the two-hop neighbours remain nodes of the same category. Recognising potential interdependencies between Service instances at each timestamp, we propose two-hop neighbour aggregation. This can be intuitively achieved by concatenating vertex features, as shown in Eq. 5 where $N_{\mathcal{S}}$ represents the number of Service instances and $S_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.
$\mathbf{X}_{\text{service}}(t)=[\mathbf{S_{1}}(t),\mathbf{S_{2}}(t),\ldots,\mathbf{S_{\mathit{N_{S}}}}(t)];$ (5)
Another method of capturing two-hop neighbours is using the Self-Attention mechanism (SA) to capture the correlations among vertices. SA works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others. This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data. See the "two hop graph structure" part in Fig. 2.
@@ 3.4 Ensemble Learning
Ensemble learning is a prevalent methodology in machine learning that reduce variance and enhance model robustness by combining multiple models. The effectiveness of ensemble learning depends on the integration of diverse, high-performing models [14]. It allows multiple models to make predictions on the same problem and vote on the final predictions. Voting can be hard or soft voting, with the mechanism such as unanimous and majority voting, which will be introduced in this subsection.
Figure 3: The ensemble learning framework used in this competition. The best-performing models: XGBoost (XGB), Support Vector Machine (SVM), and Self-Attention (SA) are combined using three voting mechanisms: hard-unanimous, hard-majority, and soft voting, to generate the prediction $\hat{y_{1}}$, where $\hat{y_{1}}\in\{0,1\}$, with 0 indicating normal and 1 indicating anomaly. Optionally, a two-stage model combination can be applied, where the unsupervised Isolation Forest further refines the anomalies detected in $\hat{y_{1}}$ to reduce false positives.
@@
As two common voting mechanisms combined with hard voting, unanimous voting classifies an instance as an anomaly only if all the models agree. Majority voting, on the other hand, classifies an instance as an anomaly if the majority of models agree.
@@ 4.1 Experimental Setup
All experiments were conducted on the ISWC 2024 ADDKG dataset, containing five days of labelled data. Three days were used for training and two for testing, with data sampled at 15 second intervals. The dataset has a significant imbalance, with anomalies representing only 4% of the total. To tackle this, class weights were incorporated into the loss functions to improve the models’ ability of anomaly detection.
As outlined in the methodology section, we tested three levels of representation. For the first representation, sequential modelling is used, where the training set is randomly divided into 80% for training and 20% for validation. For the other two representations, a more refined data splitting strategy is used. While the 80-20 split is still used for the normal class, anomalies are treated as continuous anomalous events to preserve their integrity. This ensures that the models learn from complete sequences of anomalies. For each anomaly type, one complete event is randomly selected for validation, and the remaining events are used for training.